Derive an expression for the escape energy associated with the body of mass m. Also explain the abscence of atmosphere on moon , but not on earth.

Grade:11

1 Answers

Vikas TU
7 years ago
Escape velocity at infinite distance would be defined as:
when Total Energy would be equal to zero.
That is
k.e + p.e = 0
0.5mv^2  = GMm/r
v = root(2GM/r) m/s.
The moon has no atmosphere as it has much low gravity as compare to the earth.
it is 1/6th of earth which is very very much low to bind the atomosphere of its own.